# Radiation forces and torques in optics and acoustics

### v1

@@ IV.7 Forces near interfaces
Manipulating particles near a surface is a common scenario in metamaterial-based tweezers [489], involving complex interactions between a dipole and the surrounding fields, Fig. 12. For simplicity, we consider interaction of an electric dipole with an electromagnetic wave field; this approach can be extended to magnetic or acoustic dipoles and monopoles. In the presence of an interface, the wave field at the dipole’s location can be expressed as , where is the total incident field due to the illumination and its scattering from the surface (i.e. the total field caused by the illumination and the surface if there were no particle), while is the dipole’s field back-reflected off the surface, which can be written using the Green’s function as with being the dipole’s location. The Green’s function contains all the information about the nearby environment [525, 425, 304, 416, 278, 236]. Note that does not include the primary fields radiated/scattered by the dipole (i.e. it does not include the fields from Eq. 57), which diverge at the location of the dipole itself for a point-dipole model, because the self-action of such fields is taken into account in the recoil force terms, which are left unchanged.
Two effects must be considered in this case. First, the reflected fields of the dipole induce self-polarization, resulting in a self-consistent dipole moment after infinite scatterings. Substituting here , one can define an effective polarizability [425]:
@@
Second, one must account for the field reflected from the surface as a local field modifying the force on the dipole in the general equations from Table 2: . In other words, the main force on an electric dipole near a surface (without the recoil terms) is given by [425, 416, 278, 455, 304]:
@@
The additional force term includes reflected far-field and near-field inside , introducing an array of phenomena, some of which are shown in Fig. 12(b). For example, the interaction with the reflected field can be described using the ‘image dipole’ behind the surface, which can either attract or repel the real dipole [456]. Furthermore, the near-field reflections account for surface or guided modes excited at the interface. From the momentum conservation, these surface modes result in new ‘recoil-type’ forces [565, 454, 425] (which appear from the ‘mixed’ rather than recoil term in Table 2).
@@
Other remarkable consequences of the force include levitation of the particle [456, 306, 237, 300, 455], optical binding effects, trapping and anti-trapping near metallic surfaces [278, 303], optical manipulation using hyperbolic metasurfaces [485, 277, 457, 416], magneto-optical surfaces [236], as well as novel Casimir forces [358]. The approach described here can be readily extended to the case of electric-magnetic dipoles [301] and dipole-monopole acoustic interactions.
@@ IV.8 Wavefront shaping of radiation forces and torques
In the general scattering approach to wave-induced forces and torques, Section III.1 and Fig. 5, we assume that the incident wave field is given, while the scattered field is to be found. Moreover, we assumed that the particle is surrounded by a homogeneous non-absorbing medium, so that the incident-field fluxes through a closed surface vanish. However, many setups involve a non-ideal scattering medium around the particle. In such cases, it is important to adjust the incident field to optimize it for the desired action.
@@
One of such approaches is the aberration correction to compensate the action of the surrounding medium and achieve the desired incident field (e.g., a tightly focused beam) near the particle [132, 371]. Another approach, which was introduced recently [19, 268, 411, 273], uses smart wavefront shaping outside the scattering medium and particle [129] to immediately achieve the desired action on the particle (e.g., the maximum radiation force in a given direction) rather than a specific incident field. This requires measurements of the scattering response of the system and subsequent adjustment of the incident field using this information.
@@
Since this method is universal for optical, acoustic, or other waves, it is convenient to explain it within a universal quantum-mechanical-like formalism. Assuming that the system is lossless and supports wave modes, the scattering problem is described by the unitary scattering matrix connecting the incoming and outgoing wavefields: , where the wavefunction includes all necessary wavefield components and the state vector represents the field decomposition into the modes. (Note that the ‘incoming’ and ‘outgoing’ fields are not equivalent to the ‘incident’ and ‘scattered’ fields, because part of the incident field also propagates away from the system.) The interaction of the wave with the scatterers (medium and particle) can be characterized by the Hermitian generalized Wigner-Smith operator , defined with respect to any suitable parameter of the system, [19, 268, 411]. The original Wigner-Smith operator was introduced in context of the Wigner time delay, where is the wave frequency, and the operator characterizes the time delay in the scattered field [573, 516]. Accordingly, the generalized Wigner-Smith operator describes the shift in the parameter canonically conjugated to .
For example, choosing (where is the -coordinate of the particle), the operator characterizes the shift in the momentum component of the particle, i.e., the component of the wave-induced force. Then, the largest force is achieved for the incident wavefield which is an eigenvector of with the largest eigenvalue. Obviously, to find such an optimal incident field, one has to know (measure) the scattering matrix at two neighbouring -positions of the particle. Once such measurements are done, one can choose the incident field to produce the desired -shift action on the particle, Fig. 13(a) [19, 268, 411]. In a similar manner, choosing (where is the azimuthal orientation of the particle with respect to the -axis), one can adjust the torque on the particle using the operator and its eigenvectors, Fig. 13(b) [268, 411]. Furthermore, choosing , the operator characterizes the energy shift in the system. Choosing the incident field as the eigenvector of with the minimum (most negative) eigenvalue allows one to achieve optimal cooling in a multiple-particle system [273].
@@
This smart wavefront shaping of radiation forces and torques has been demonstrated both for electromagnetic [268] and acoustic [411] fields. It was also applied for achieving the optimal pulling force [269] and improving the stiffness of optical tweezers [122]. The advantages of this approach are evident: its universality and optimized action for any scattering medium and arbitrary-shaped particles. The drawbacks are: the lossless character of the system (to deal with Hermitian operators ) and time-consuming measurements of the full scattering matrix at each step of the system evolution.
@@ IV.9 Photophoretic and acoustophoretic forces
The medium surrounding the particle can also affect the particle’s dynamics via thermal effects. A non-uniform absorption of the radiation produces temperature gradients in the particle and surrounding medium, which in turn yields radiation-induced flows and photophoretic forces [194, 595, 458, 534, 270, 164].
Studies of photophoretic forces have been primarily focused on atmospheric aerosol particles [546, 448, 299, 575, 146, 459, 176, 145, 297], but recently they significantly extended the range of potential applications, including optical manipulation of nanocasters [496], ‘giant’ () particles [498], chiral nanotube sorting [515], controllable pulling forces [355], volumetric displays [514], levitation and near-space flights [158, 47], etc.
In acoustics, there are acoustophoretic forces which are induced by the interplay of the sound wave field and properties of the surrounding medium, and which play a crucial role in acoustofluidics [328, 343, 116, 117, 296]. However, the acoustofluidic literature often calls all sound-wave-induced forces, including the acoustic forces described in the previous Sections, as “acoustophoretic”. In the framework of our review, it makes sense to separate different mechanism and refer to acoustophoretic forces as those which are not described by the linear acoustic-scattering theory. Such forces can appear due to nonlinear hydrodynamical, viscous, or thermal effects in the surrounding medium.
@@ IV.10 Ponderomotive forces on the medium particles
So far, we have considered the interaction of an incident wave field with a particle immersed in a homogeneous optical or acoustic medium. A wave propagating in a medium also acts on the microscopic particles (molecules) constituting this medium. To describe these effects, one has to analyse the microscopic equations of motion governing the medium’s particles.
We start with a sound wave propagating in a fluid or gas. To describe the most interesting force effects, let us consider a quasi-monochromatic wave with a central frequency , which can change its amplitude not only in space but also in time (e.g., a propagating wavepacket). Then, the cycle-averaged force density acting on the fluid particles can be written as [93]:
Here the first term is a ponderomotive gradient force, originally derived for a charged particle in an electric wave field [222], while the second term is related to the temporal change in the local canonical momentum density of the wave field. Note the similarity between Eq. (46) and the dipole acoustic force on a particle, Eq. (22). The gradient force tends to expel the medium particles from the zones of high kinetic wave energy density.
Considering a wave front coming to a given point, starting from zero amplitude and saturating at a constant amplitude, the second term in Eq. (46) will accelerate the medium particles to a constant velocity given by
This is the Stokes drift known in fluid mechanics [113, 103, 93, 520], and providing the microscopic mechanical origin of the acoustic canonical momentum density , Fig. 14. Thus, the ponderomotive force (46) is closely related to the origin of the canonical wave momentum.
Let us consider now propagation of an electromagnetic wave in a plasma-like medium (e.g., a metal) with freely moving electrons. Remarkably, the ponderomotive force acting on the medium’s electrons is given exactly by the first line in Eq. (46), where and are the velocity and density of the electrons [93]. The gradient and time-derivative parts of this force are related to the electron contributions to the electromagnetic wave energy and canonical momentum in the dispersive medium [93, 88, 90]. The medium dispersion, , inevitably appears from the microscopic interaction of the medium electrons with the electromagnetic field. In the presence of such dispersion, the electromagnetic energy and canonical momentum equations (2) and (4) are modified by the substitution [316, 88, 90]. In a dielectric medium, similar conclusions can be derived by considering the interaction of an electromagnetic wave field with polarizable atoms behaving as electric dipole particles [380, 243].
@@ V Examples of forces and torques in structured fields
In this section we describe the most important examples of optical and acoustic forces and torques that appear in specific systems and applications.
@@ V.1 Evanescent waves
We start with a simple example exhibiting nontrivial forces and torques, which is important as a useful model and also relevant to numerous wave systems. Namely, we consider a particle immersed in a single evanescent wave [209], i.e., a wave propagating along the -axis and decaying in the -direction: , where . Such wave can be regarded as a plane wave with the complex wavevector . Here and hereafter the overbars denote the unit vectors of the corresponding axes.
@@ V.1.1 Optics
A generic electromagnetic evanescent wave field can be written as [87, 100]
@@
Here is a complex number determining the polarization state of the wave, such that , , and are the three normalized Stokes parameters () describing, in the limiting plane-wave case , the degrees of the vertical/horizontal, diagonal, and right-hand/left-hand circular polarizations, respectively. Substituting the fields (48) into Eqs. (2)–(6), we obtain the distributions of the energy, momentum, and spin densities:
@@
Note also that the helicity density (6) and the “chiral momentum” density (35) are determined by the parameter : and . Substituting Eqs. (49) into Eqs. (22) and (23) or (26), we find that an isotropic Rayleigh particle in an evanescent electromagnetic field experiences: (i) an -directed gradient force due to the energy-density (intensity) gradient, (ii) a -directed radiation-pressure force due to the canonical momentum (phase gradient), and (iii) a radiation torque due to the spin density, which generally has all three components, see Fig. 15. In the simplest case of a vertically polarized wave (, ), an absorbing electric-dipole particle will experience a transverse -directed torque due to the transverse spin , Fig. 15(c,d). It is worth remarking that such helicity-independent spin has been recognized rather recently [99, 87, 100, 10, 395, 195], and it has found exciting applications for efficient spin-momentum coupling using evanescent and guided waves, both in optics [453, 422, 204, 105, 104, 352] and acoustics [488, 102, 589, 606, 599, 567]. Thus, the geometry of the main optical forces and torque in a vertically-polarized evanescent wave is entirely similar to the one shown schematically in Fig. 7.
@@
In addition to the main forces and torque, the evanescent wave produces a nontrivial recoil force (24). Indeed, the complex kinetic momentum density for an evanescent field (48) equals
@@
Here the -dependent -directed term corresponds to the transverse circular-polarization-dependent component of the Poynting vector first noticed by Fedorov and Imbert [206, 205, 274]. Since the complex kinetic momentum (50) generally has a nonzero component, it can produce a transverse recoil force (24) orthogonal to both the gradient and radiation-pressure forces. Such lateral recoil force was described theoretically in [87] and measured experimentally in [25, 345], Fig. 15(a,b).
@@
The presence of transverse spin suggests that an evanescent wave can produce a transverse chiral force on chiral particles. The main terms in the chiral force (IV.2.1) produce an -directed helicity-gradient force and a -directed force due to the “chiral momentum”. In turn, the recoil terms in Eq. (IV.2.1) have a nonzero -component due to the transverse spin. Such chiral recoil force, suitable for chiral sorting of particles, was described in [565, 257, 366, 239].
@@
Note that since evanescent waves appear near interfaces (e.g., upon total internal reflection or in surface waves), particles in an evanescent field are inevitably located close to an interface. Therefore, the interface-induced forces, Section IV.7, can also play an important role in such systems.

### v3

@@ IV.7 Forces near interfaces
When a particle is placed near an interface or another object, the force it experiences is no longer limited to the principal forces described in Section III. The presence of the surface introduces additional contributions coming from the back-reflected scattered fields, Fig. 12. Manipulating particles near a surface is a common scenario in plasmonic and metamaterial-based tweezers [684, 541]. For simplicity, we focus on the interaction of an electric dipole with an electromagnetic wave field; this approach extends to magnetic or acoustic dipoles and monopoles.
@@
In the presence of an interface, the total wave field at the dipole’s location can be written as , where is the total incident field due to the illumination and its scattering from the surface (i.e. the field that would exist without the particle), and is the dipole’s own field reflected back from the surface. The latter can be expressed using the Green’s function: with denoting the dipole’s position. The Green’s function encodes all information about the surrounding environment [580, 464, 325, 455, 295, 249, 599, 322]. Note that does not include the primary fields radiated/scattered by the dipole, Eq. (59), which diverge at the dipole’s location in the point-dipole model; the self-action of such fields is taken into account in the recoil force terms, which remain unchanged.
Two effects must be taken into account near a surface. First, the reflected dipole field induces self-polarization, leading to a self-consistent dipole moment , which includes all orders of multiple scattering. Substituting , yields an effective polarizability [464, 599]:
@@
Second, The reflected field also acts directly as a local field that modifies the force on the dipole in the general equations of Table 2: . As a result, the main force on an electric dipole near a surface (without recoil terms) is given by [464, 455, 295, 500, 325]:
@@
The additional force term includes both the reflected far-field and the near-field contained in , giving rise to a variety of phenomena, some of which are illustrated in Fig. 12(b). For instance, the interaction with the reflected field can be interpreted in terms of an ‘image dipole’ behind the surface, which may either attract or repel the real dipole [501]. Furthermore, the near-field reflections account for the excitation of surface or guided modes at the interface. By momentum conservation, these modes lead to new ‘recoil-type’ forces [625, 499, 464, 366, 402] (which originate from the ‘mixed’ rather than recoil term in Table 2).
@@
Other remarkable consequences of the force include particle levitation, [501, 327, 250, 319, 500], optical binding, trapping or anti-trapping near metallic surfaces and nanostructures [310, 295, 324, 403, 684], optical manipulation using metamaterials [541], hyperbolic metasurfaces [535, 294, 502, 455], magneto-optical surfaces [249], and novel Casimir forces [390]. For instance, particle levitation is produced by the phase-dependent interaction between the source dipole and its image; by appropriately tuning the reflection coefficients, one can induce a repulsive regime in which the dipole and its image effectively repel each other. In turn, Casimir forces for an electric dipole near a planar surface arise from stochastic fluctuations of both the dipole moment and the field generated by currents in the surface, both of which result in a stochastic contribution to the force . The framework described here can also be extended to electric-magnetic dipoles [320] and to dipole-monopole acoustic interactions.
@@ IV.8 Wavefront shaping of radiation forces and torques
In the general scattering approach to wave-induced forces and torques, Section III.1 and Fig. 5, we assume that the incident wave field is given, while the scattered field is to be determined. Moreover, we have assumed that the particle is embedded in a homogeneous lossless medium, so that the incident-field fluxes through a closed surface vanish. In practice, however, many setups involve a non-ideal surrounding medium. In such cases, it becomes important to adjust the incident field to optimize the desired effect on the particle.
@@
One strategy is aberration correction, which compensates distortions introduced by the medium in order to reproduce a target incident field (e.g., a tightly focused beam) at the particle’s location [139, 405]. Another approach, which was introduced recently [21, 283, 448, 289], employs smart wavefront shaping outside the scattering medium and particle [136] to directly optimize the desired action on the particle (e.g., the radiation force in a given direction) rather than a specific incident field. This method requires measuring the system’s scattering properties and then adapting the illumination accordingly.
@@
Since this method is universal for optical, acoustic, or other waves, it is convenient to explain it within a universal quantum-mechanical-like formalism. Assuming that the system is lossless and supports wave modes, the scattering problem is described by a unitary scattering matrix that connects incoming and outgoing waves: , where the wavefunction includes all necessary field components and the state vector represents the field decomposition into the modes. (Here ‘incoming’ and ‘outgoing’ refer to modal propagation directions and should not be confused with the ‘incident’ and ‘scattered’ fields, since part of the incident field also contributes to outgoing channels.) The interaction of the wave with the scatterers (medium and particle) can be characterized using the Hermitian generalized Wigner-Smith operator , defined with respect to any suitable parameter of the system, [21, 283, 448, 102]. The original Wigner-Smith operator was introduced in the context of the Wigner time delay, where is the frequency, and describes the temporal delay of the scattered field [633, 570]. More generally, the operator quantifies the shift in the variable canonically conjugate to .
For example, choosing (the -coordinate of the particle), the operator characterizes the shift in the momentum component of the particle, i.e., the component of the wave-induced force. The maximum force is achieved when the incoming field is the eigenvector of corresponding to the largest eigenvalue. To find this optimal incoming field, one has to determine the scattering matrix for two neighboring -positions of the particle (to calculate its -derivative in the operator ). Once the operator is known, the incident field can be tailored to realize the desired force on the particle, Fig. 13(a) [21, 283, 448].
In a similar manner, choosing (the azimuthal orientation of the particle with respect to the -axis) allows one to control the torque using the operator and its eigenvectors, Fig. 13(b) [283, 448]. Furthermore, choosing , the operator quantifies energy shifts in the system. Choosing the incident field as the eigenvector of associated with the most negative eigenvalue enables optimal cooling in a multi-particle system [289].
@@
This smart wavefront shaping of radiation forces and torques has been demonstrated in both electromagnetic [283] and acoustic [448] systems. It has also been applied to realize optimal pulling forces [284] and to enhance the stiffness of optical tweezers [128]. The advantages of this method are its universality and ability to optimize action for particles of any shape in arbitrary scattering environments. Its limitations are the assumption of a lossless system (needed to ensure that is Hermitian) and time-consuming measurements of the full scattering matrix at each step of system evolution.
@@ IV.9 Photophoretic and acoustophoretic forces
The medium surrounding a particle can also influence its dynamics via thermal effects. Non-uniform absorption of radiation produces temperature gradients within the particle and surrounding medium, which in turn generate radiation-induced flows and photophoretic forces [204, 657, 503, 589, 285, 173].
Research on photophoretic forces has traditionally focused on atmospheric aerosol particles [605, 489, 318, 635, 154, 504, 185, 153, 316]. More recently, however, the scope has broadened to include applications such as optical manipulation of nanoclusters [549] and ‘giant’ particles of order [551], chiral nanotube sorting [569], controllable pulling forces [386], volumetric displays [568], and levitation and near-space flights [167, 50].
In acoustics, analogous acoustophoretic forces arise from the interplay between sound waves and the properties of the surrounding medium, and they play a crucial role in acoustofluidics [350, 373, 122, 123, 315]. However, in the acoustofluidics literature, the term “acoustophoretic” is often used broadly to describe all sound-wave-induced forces, including those covered in previous sections. Within the framework of this review, it is useful to distinguish mechanisms: we reserve the term acoustophoretic forces for those not captured by linear acoustic-scattering theory, such as forces arising from nonlinear hydrodynamic, viscous, or thermal effects in the medium.
@@ IV.10 Ponderomotive forces on the medium particles
So far, we have considered the interaction of an incident wave field with a particle immersed in a homogeneous optical or acoustic medium. However, a propagating wave also acts on the microscopic constituents of the medium itself (e.g., molecules or fluid elements). To describe such effects, one has to analyze the microscopic equations of motion governing the medium’s particles.
Let us consider a sound wave propagating in a fluid or gas. To describe the most relevant force effects, we consider a quasi-monochromatic wave with central frequency , whose amplitude may vary both in space and in time (e.g., a propagating wavepacket). The corresponding cycle-averaged force density acting on the fluid particles can be expressed as [97, 87]:
Here the first term represents a ponderomotive gradient force, originally derived for a charged particle in an oscillating electromagnetic field [235], while the second term corresponds to the temporal variation of the local canonical momentum density in the wave field. Note the similarity between Eq. (48) and the dipole acoustic force on a particle, Eq. (24). The gradient force tends to expel the medium’s particles from regions of high kinetic wave energy density.
Consider a wavefront arriving at a given point, starting from zero amplitude and saturating at a constant value. In this case, the second term in Eq. (48) accelerates the medium’s particles to a steady velocity
This is the Stokes drift known in fluid mechanics [118, 108, 574, 87], which provides the microscopic mechanical origin of the acoustic canonical momentum density , Fig. 14. Thus, the ponderomotive force (48) is intimately linked to the origin of canonical wave momentum.
Now consider the propagation of an electromagnetic wave in a plasma-like medium (e.g., a metal) with freely moving electrons. Remarkably, the ponderomotive force acting on the electrons is given by the first line in Eq. (48), with and representing the electron velocity and density [97]. The gradient and time-derivative parts of this force correspond to the electron contributions to the electromagnetic wave energy and canonical momentum in a dispersive medium [97, 92, 94]. The medium dispersion, expressed through , arises from the microscopic electron-field interaction. In this case, the electromagnetic energy and canonical momentum equations (2) and (6) are modified by the substitution [337, 92, 94]. In dielectric media, analogous results can be derived by considering the interaction of an electromagnetic wave field with polarizable atoms behaving as electric dipoles [415, 256].
@@ V Examples of Forces and Torques in Structured Fields
In this section, we describe key examples of optical and acoustic forces and torques that arise in specific systems and applications.
@@ V.1 Evanescent waves
We begin with a simple yet illustrative example that exhibits nontrivial forces and torques. This case is important both as a model system and because of its relevance to numerous real systems. Namely, we consider a particle immersed in a single evanescent wave [222], i.e., a wave propagating along the -axis and decaying in the -direction: , where . Such a wave can be viewed as a plane wave with a complex wavevector . Here and hereafter, the overbars denote unit vectors along the corresponding axes.
@@ V.1.1 Optics
An electromagnetic evanescent wave field can be written as [91, 105]
@@
Here is a complex number that determines the polarization state of the wave. The corresponding normalized Stokes parameters are: , , and , with . In the limiting plane-wave case (), , , and describe the degrees of vertical/horizontal, diagonal (), and right-hand/left-hand circular polarizations, respectively.
Substituting the fields (50) into Eqs. (2)–(8), we obtain the the energy, momentum, and spin densities:
The helicity density (8) and the “chiral momentum” density (37) are determined by the parameter : and .
Substituting Eqs. (51) into Eqs. (24) and (25) or (28), we find that an isotropic Rayleigh particle in an evanescent electromagnetic field experiences: (i) an -directed gradient force due to the energy-density (intensity) gradient, (ii) a -directed radiation-pressure force due to the canonical momentum (phase gradient), and (iii) a radiation torque due to the spin density, which generally has all three components, see Fig. 15.
@@
In the simplest case of a vertically polarized wave (, ), an absorbing electric-dipole particle experiences a transverse -directed torque caused by the transverse spin , Fig. 15(c,d). Notably, this helicity-independent spin component was recognized only recently [104, 91, 105, 11, 432, 205]. It has since found applications for efficient spin-momentum coupling using evanescent and guided waves, both in optics [498, 461, 215, 110, 109, 383] and acoustics [540, 107, 651, 671, 661, 627]. Note that the geometry of the main optical forces and torque in a vertically-polarized evanescent wave is entirely similar to that shown schematically in Fig. 7.
@@
In addition to the main forces and torque, the evanescent wave produces a nontrivial recoil force (26). The complex kinetic momentum density of the evanescent field (50) is
@@
Here, the -dependent -directed term corresponds to the transverse, circular-polarization-dependent component of the Poynting vector first identified by Fedorov and Imbert [217, 216, 291]. The component of the complex kinetic momentum (52) can produce a transverse recoil force (26) orthogonal to both the gradient and radiation-pressure forces. This lateral recoil force was described theoretically in [91] and measured experimentally in [27, 375], Fig. 15(a,b).
@@
Furthermore, the presence of transverse spin suggests that an evanescent wave can exert a transverse chiral force on chiral particles. The leading terms in the chiral force (IV.2.1) yield an -directed helicity-gradient force and a -directed force associated with the “chiral momentum”. In turn, the recoil terms in Eq. (IV.2.1) have a nonzero -component due to the transverse spin. This chiral recoil force, suitable for chiral particle sorting, has been analyzed in [625, 271, 398, 252].
@@
Note that since evanescent waves appear near interfaces (e.g., under total internal reflection or in surface waves), particles in such fields are located close to an interface. Therefore, the interface-induced forces, described in Section IV.7, can also play an important role in these systems.